OpenStudy (anonymous):

solve; 10ac-x/11=-3, for a

OpenStudy (anonymous):

10ac-x/11=-3 -10ac -10ac get a on one side - seperate from x -x/11 = -3 -10ac +3 +3 add 3 to both sides -x/11 +3 = -10ac divide everything by -10c to get a alone x/110c +3 = a

OpenStudy (anonymous):

Thank you so much